OpenStudy (anonymous):

Solve the system x + y = 6 and x - y = 2. What is the value of x? A. 6 B. 2 C. 4 D. 0

OpenStudy (anonymous):

C. Add both equations tgt. (x+y)+ (x-y) = 6+2 2x = 8 x=4

OpenStudy (anonymous):

yep c